# Detecting Harmful Memes with Decoupled Understanding and Guided CoT Reasoning

### v1

@@ 4 Experiments
Table 1 shows that both Qwen2.5-14B and Mistral-12B outperform the two retrieval-augmented baselines on FHM and HarMeme, even surpassing LoReHM’s performance based on the more advanced GPT-4o. Specifically, Qwen2.5-14B can further outperform LMM-RGCL+RKC on MultiOFF, MAMI and PrideMM. Compared to low-resource baselines without retrieval augmentation, small LLMs under our guided zero-shot CoT prompting can also significantly outperform both LMM-RGCL and Mod-HATE that are evaluated under few-shot settings. We report the zero-shot performance of GPT-4o-mini based on direct multimodal inference, while the performance of GPT-4o is sourced from Huang et al., 2024. As shown, small LLMs under U-CoT+ consistently outperform or match GPT-4o(-mini) on five out of six datasets in Table 1, although showing slightly weaker performance on PrideMM. Table 2 compares the best zero-shot results achieved by Qwen2.5-14B and Mistral-12B models to those by GPT-4V on Goat-Bench Lin et al. (2024b) (See Table 7 for complete results). The performance of GPT-4V is based on multimodal zero-shot CoT prompting (M-CoT), while we report our best results both with and without providing guidelines to the LLMs. As can be seen, small LLMs achieve highly comparable performance to, and in some cases even outperform, GPT-4V across all four sub-tasks.
These results highlights the advantage of our method in terms of resource efficiency and flexibility, especially when compared to approaches based on supervised learning or more advanced multimodal models. It demonstrates the inherent capabilities of LLMs to detect harmful memes across diverse domains and contexts in zero-shot settings by effectively following explicit guidelines, whereas supervised models often struggle with poor out-of-distribution generalization. Despite results given by pretrained LLMs with more than 10 billion parameters, we do not observe equally competitive performance on smaller LLMs (e.g., Qwen2.5-7B and Llama3.1-8B). This showcases that harmful meme classification in a text-only unimodal setting still places certain demands on the overall reasoning capabilities of the underlying LLMs.
4.3 Ablation Study
We present the results of ablation studies in Table 3 and Table 6, focusing on the performance comparison of different prompting schemes: M-CoT, U-CoT, and U-CoT+. Figure 2 presents the detailed confusion matrices under the three different zero-shot CoT schemes. To ensure fair comparisons between multimodal and unimodal settings, we only replace the input images of memes with our high-fidelity meme descriptions while keeping all textual prompts identical across the compared schemes. As shown, under M-CoT, the 7B Qwen2VL and Llava1.6 models perform similarly on HarMeme, Harm-P, and MAMI, while Qwen2VL outperforms Llava1.6 on the other three datasets. As shown in Figure 2, except for PrideMM, the two LMMs exhibit an inherent bias toward the negative class across most datasets, frequently overclassifying memes as “harmless”. This reflects the limited capability of small LMMs in detecting harmfulness in memes based solely on visual inputs. Converting the default multimodal settings into unimodal, text-only settings using more capable LLMs can, in most cases, improve meme classification performance by balancing predictions, leading to more examples being classified as “harmful”. Consistent improvements from M-CoT to U-CoT are observed on FHM, Harm-P, and MAMI. However, on HarMeme and MultiOFF, only meme descriptions based on Qwen2VL-7B helps Qwen2.5-14B better detect “harmful” memes. Comparing U-CoT to U-CoT+ demonstrates the overall effectiveness of our human-crafted guidelines, which is particularly prominent for HarMeme, MAMI and PrideMM. For example, our human-crafted guidelines tend to make Qwen2.5-14B produce more predictions of “harmful” on FHM and more predictions of “harmless” on PrideMM, but have limited impact on its classification decisions for the two political meme datasets, MultiOFF and Harm-P.
Huang et al., 2024 presents GPT-4o-generated insights for harmful meme detection, which serve a similar purpose to our human-crafted guidelines by providing additional instructions to guide model reasoning. To evaluate their effectiveness relative to ours, we conduct an ablation study in which we replace our human-crafted guidelines with these GPT-4o-generated insights (results shown in Table 4). We observe that GPT-4o-generated insights do not significantly improve the classification performance of small LLMs and may even lead to performance drops when applied to Mistral-12B. This suggests that small LLMs fail to fully leverage GPT-generated insights, which are often too general and verbose and lack clear context-specific relevance and distinctions between individual rules. This further highlights the effectiveness of our human-crafted guidelines.
Figure 2: Comparison of classification confusion matrices under three different zero-shot CoT prompting schemes, with Llava1.6-7B and Qwen2VL-7B as the base LMMs for Meme2Text, and Qwen2.5-14B as the base LLM for unimodal, text-only inference.
4.4 Error Analysis
@@
Our method explicitly provides detailed rationales from which predicted labels are derived, and this transparency in the decision-making process allows us to analyze the inherent behaviors of LLMs for the task of harmful meme detection. We randomly sample 30 balanced error cases (False Negative:False Positive = 1:1) generated by Qwen2.5-14B under U-CoT+ from each of the three relatively underperformed datasets: FHM, Harm-P, and PrideMM. We identify and summarize the common causes of misclassifications by closely examining the step-by-step reasoning outputs (see Appendix D for detailed case study): (i) Incorrect or missing visual details: Due to limited image understanding and text recognition capabilities, 7B LMMs often struggle to generate accurate details for visually noisy, text-heavy memes such as edited screenshots of chats or tweets. They also show limited robustness in recognizing celebrities, including historically significant figures such as Anne Frank. (i) Excessive censorship: LLMs tend to overinterpret meme content or overstate the potential negative effects when they interpret and apply the guidelines too rigidly or adopt an overly strict perspective, such as regarding sarcasms or parodies on political debates as trivializing serious issues. (i) Misinterpretation: On one hand, LLMs may misinterpret the meme’s tone or intent, either in an opposite or skewed direction or fail to grasp the intended meaning when the meme is subtle, ambiguous or poorly framed. On the other hand, discrepancies may exist between LLMs’ and human interpretations of the guidelines, leading the models to overlook or fail to apply the guidelines consistently. (i) Different interpretations: LLMs may arrive at differing yet reasonable interpretations, often due to discrepancies between the LLMs’ and annotators’ perceptions of the degree of harm. This means that LLMs’ classifications are not necessarily incorrect, even when they deviate from annotator labels. For instance, LLMs often classify some examples annotated as “harmful” as “harmless”, interpreting them as merely mild ridicule or satire, which is both common and acceptable in political memes. (i) Misclassification of targets: Since the predicted target determines which subset of guidelines is applied, inaccurate predictions of the meme’s target subject can lead to guideline mismatches and, consequently, to misclassifications. (i) Blind Spot: LLMs can still be limited by types of harmfulness not covered by the guidelines, particularly when encountering rare examples that deviate from the main context.
5 Conclusion

### v2

@@ 4.1. Comparing U-CoT+ to Baselines
U-CoT+ vs. Low-Resource Baselines Low-resource baselines with retrieval-based augmentation such as LoReHM, MInD, though operating under a zero-shot setting, further benefit from a majority voting of ground-truths of the top-K most similar memes retrieved from the embedding space. In contrast, U-CoT+ operates under a more constrained low-resource setting, without additional access to auxiliary ground-truth signals from neighbour memes as reference for decision makings. Despite the absence of retrieval augmentation, we show that lightweight LLMs can unlock their full potential for zero-shot harmful meme detection. As shown in Table 1, with the support of our human-crafted guidelines, both Qwen2.5-14B and Mistral-12B can largely outperform MInD, and can outperform the other two retrieval-augmented baselines on FHM and HarMeme, even surpassing LoReHM’s performance based on GPT-4o. Qwen2.5-14B can further outperform LMM-RGCL+RKC on MultiOFF, MAMI and PrideMM. Notably, the diminished out-of-distribution performance of LMM-RGCL, even when enhanced by the retrieval mechanism (RKC), further demonstrates the limited generalizability of SFT methods. Furthermore, U-CoT+ can also significantly outperform LMM-RGCL+FS and Mod-HATE, both of which are baselines without retrieval augmentation but are evaluated under few-shot settings. Besides, lightweight LLMs under U-CoT+ consistently outperform or match GPT-4o-mini’s zero-shot performance based on multimodal inference across five of six datasets, with slightly weaker results only observed on PrideMM. Table 2 compares the zero-shot performance of our method to that of GPT-4V on Goat-Bench (Lin et al., 2024b). As can be seen, U-CoT+ enables lightweight unimodal LLMs to achieve performance highly comparable to, and in some cases even surpassing that of the advanced multimodal GPT-4V across all four sub-tasks.
In summary, results validate U-CoT+’s advantages: (i) It effectively leverages lightweight LMMs and LLMs, and achieves highly comparable zero-shot performance relative to more resource-intensive baselines, thereby demonstrating strong resource efficiency. (ii) Its underlying pipeline and principles support empirical, context-specific adaptation to diverse detection criteria with only minor human efforts, offering flexibility. (iii) Its consistently top-ranked performance among low-resource methods across diverse datasets underscores the scalability of our guideline crafting process, which generalizes well across varied sociocultural contexts.
Figure 2. Confusion matrices of Qwen2.5-14B based on meme descriptions sourced from different 7B LMMs.
4.2. Ablation Study
Table 3 presents the main results of ablation studies that compare different prompting schemes and validate the effectiveness of key framework components. Figure 2 presents the detailed confusion matrices by Qwen2.5-14B.
Unimodal vs. Multimodal. We observe consistently improved zero-shot performance on FHM, Harm-P, and MAMI merely by converting the default multimodal setting to unimodal, text-based inference. This highlights the effectiveness of modality conversion, which enables certain meme messages to be more explicitly revealed when conveyed in text. Figure 2 shows that both 7B LMMs exhibit an inherent bias toward the negative class, frequently overclassifying memes as “harmless” in most datasets. This evidences their limitation in detecting harmfulness based on visual inputs. Reasoning with more capable unimodal LLMs can, in most cases, help balance classification with more memes revealed to be “harmful”.
Effectiveness & Robustness. Comparing U-CoT+ against U-CoT results demonstrates the overall effectiveness of our human-crafted guidelines. Notably, U-CoT+ enables LLMs of comparable size to 7B LMMs to outperform M-CoT results across four datasets. We further apply different perturbation strategies e.g., Rephrasing, Shuffling, and Modifying to our guidelines and compute the resulting average performance gain of U-CoT+ over U-CoT. As shown by $\bar{\Delta}_{\textit{U-CoT+}}$ in Table 3, despite the potential randomness, semantic nuances and logic flow disruption introduced, our guidelines remain conceptually constructive, consistently improving performance across most datasets. This also evidences U-CoT+’s general robustness to guideline perturbations.
Human-crafted vs. GPT-generated. We compare the effectiveness of our human-crafted guidelines with GPT-4o-generated harmful meme detection insights from LoReHM (Huang et al., 2024). As show in Table 3 ($\bar{\Delta}_{\textit{U-CoT+}}$ vs. $\Delta_{\textit{U-CoT+L}}$), prompting LLMs with human-crafted guidelines in U-CoT+ yields greater improvements over zero-shot U-CoT than using GPT-generated insights, which may even lead to performance drops when applied to Mistral-12B. These GPT-4o-summarized insights alone, which tend to be overly general and verbose, lacking explicit context-specific relevance and rule-level diversity, prove insufficient for effectively guiding lightweight unimodal LLMs in harmful meme detection. This further underscores the usability and effectiveness of our guideline crafting principles in providing targeted, interpretable guidance that is easy to follow.
Guidelines vs. Few shots. We evaluate four few-shot (FS) settings (4, 6, 8 and 10-shot) with class-balanced meme descriptions randomly sampled from the holdout pool (see Section C.4 for details of demonstration example construction.), and compare their average performance gain over zero-shot U-CoT ($\bar{\Delta}_{\textit{U-CoT+FS}}$) to $\bar{\Delta}_{\textit{U-CoT+}}$ in Table 3. Results show that, despite occasional improvements, providing LLMs with meme descriptions as few-shot demonstration examples is generally less effective than guideline-based prompting, and may even result in deteriorated performance compared to that of zero-shot inferences, suggesting that LLMs’ in-context learning ability does not translate well to this task, which aligns with previous findings for LMMs (Mei et al., 2025; Huang et al., 2024).
7B LMMs vs. GPT-4o-mini for Meme2Text. We replace meme descriptions produced by our proposed pipelines with GPT-4o-mini-generated descriptions, which generically provide more accurate visual information. Results for U-CoT+ in Table 3 show that reasoning over meme descriptions generated by the more capable GPT-4o-mini does not consistently outperform those produced by less advanced lighter-weight LMMs, suggesting that the performance bottleneck is not primarily visual. This further indicates that our proposed meme2text pipeline, despite relying on low-resource models, is still effective in capturing visual information essential for harmfulness assessment.
Single vs. Ensemble. We compare single-LMM U-CoT+ with ensemble strategies that aggregate predictions from other scheme variants (i.e., M-CoT and U-CoT), as well as from U-CoT+ based on different LMMs (See Table 6). Results in Table 3 show that even without relying only on a single best-performing 7B LMM, we can largely preserve performance through appropriate ensembles of predictions, which may even yield further improved results on MAMI and PrideMM. In practice, ensemble strategies help mitigate uncertainty in LLMs’ sensitivity to the choice of 7B LMMs for visual information extraction, while also alleviating potential over-censorship introduced by strict guideline-based assessment.
Figure 3. Comparing LLM rationales before and after applying context-specific guidelines. Incorrect. Correct.
Figure 4. Examples of incorrectly classified memes and their corresponding error type.
5. Case Study
Although grounded in a low-resource setup, U-CoT+ fully leverages the inherent knowledge and reasoning capabilities of lightweight LLMs for explainable harmful meme detection by prompting them to generate explicit step-by-step rationales leading to the final predictions. Comparing LLMs’ reasoning trajectories enables a closer examination of how our human-crafted guidelines influence their “mindset”, as illustrated in Figure 3. First, the guidelines provide LLMs with more aligned and context-specific guidance that enhances their associations, particularly for those negative interpretations with implicit harmful cues. For example, in Figure 3(a), when no guideline is provided, despite correctly recognizing the visual content, the LLM misinterprets the mention of “plane” as “owning a plane”, linking it to a neutral stereotype that assumes Middle Eastern individuals as wealthy. In contrast, when guidelines are presented, the LLM correctly identifies the sensitivity of the “plane” reference and associates it with a more serious implication of terrorism by pointing out its possible allusion to the 9/11 attacks, which is crucial for determining the meme as harmful. Second, the guidelines clarify the boundary between black and white by defining detailed checklists, thereby reducing uncertainty in the “gray zone”. For the meme shown in Figure 3(b), when guidelines are absent, the LLM reasonably acknowledges the meme’s potential offensiveness but ultimately downplays it as common political satire. With the guidelines in place, however, the LLM becomes more confident in affirming the meme as offensive based on matches with multiple criteria, such as misinformation, stereotyping, and personal attacks etc.  As such, U-CoT+ enables human content moderators to intuitively formulate and apply detection guidelines, and to further refine them or verify models’ verdicts based on reasoning outputs, thereby supporting more consistent and accountable decisions.
6. Error Analysis
@@
U-CoT+’s transparency in decision makings enables fine-grained error analysis that reveals the inherent behaviors of lightweight LLMs (and LMMs) in harmful meme detection. We summarize common causes of misclassification by closely examining LLMs’ rationales. Figure 4 presents typical misclassified examples and Figure 6 shows the quantitative distribution of different error types: (i) Incorrect or missing visual details: The image understanding and text recognition capabilities of 7B LMMs remain limited. For instance, the LMM fails to recognize Anne Frank in Figure 4(a), thereby missing the connection between the visual element “ash” and the insensitive caption. Moreover, LMMs still struggle to provide accurate details for visually noisy, stylistically abstract, blurred, or text-heavy memes, such as edited screenshots of chats or tweets. (ii) Excessive censorship: LLMs tend to overinterpret or overstate potential negative effects when they apply the guidelines too rigidly or adopt an overly strict perspective, such as interpreting sarcasm or parody in political debates shown in Figure 4(b) as trivializing serious issues. (iii) Misinterpretation: LLMs may misinterpret a meme’s tone or intent when the meme is subtle, implicit, ambiguous, or poorly framed. Discrepancies between LLMs’ and human interpretations of the guidelines can also cause LLMs to overlook or inconsistently apply them. (iv) Different interpretations: LLMs may arrive at reasonable interpretations that diverge from human annotations, largely due to differences in perceived levels of harm. In such cases, LLM judgments are not necessarily incorrect. For instance, LLMs may classify some annotated harmful memes as harmless by interpreting them as mild ridicule or satire that is common and generally acceptable in political discourse. These misclassifications may instead reflect inconsistent or erroneous annotations. (v) Target mismatch: The inferred target subject of a meme determines which subset of guidelines is applied. Discrepancy in target identification can lead to guideline mismatches and subsequent misclassifications. Figure 4(e) is labeled as harmful toward President Biden, but the LLM instead interprets the target as the LGBTQ+ community and thus predicts the meme as harmless. (vi) Blind Spot: LLMs adhere to guideline-bounded reasoning and do not actively extend beyond it to account for long-tailed harmfulness types outside the main context. In Figure 4(f), the LLM fails to generalize to the vegan group under the PrideMM guidelines, which primarily focus on LGBTQ+ issues. These errors highlight limitations in both models and datasets, motivating future improvements: (i) Enabling general LLMs to rapidly learn and adapt to evolving popular culture and communication practices; and (ii) Developing multi-level, fine-grained, caption- and explanation-aware annotation frameworks for high-quality harmful meme detection benchmarks.
7. Conclusion